PLEASE HELP! THANK YOU.
Solve my substitution method.
6x+5 = 5 and
x = 30 - 5y. I know that in the second equation x is already solved for so I would substitute this into equation one right?
6(30 - 5y) + 5 = 5; 180 - 30y +5 = 5;
NOT SURE AFTER THIS.

Answer by nerdybill(7384)

You have a typo for your first equation:
6x+5 = 5
It is missing a 'y' somewhere...
We need this to answer the question
.
6x+5y = 5
x = 30 - 5y
.
6x + 5y = 5
x + 5y = 30
.
Multiply both sides of bottom equation by -1 and add to top equation:
6x + 5y = 5
-x - 5y = -30
----------------
5x = -25
x = -5
.
substitute above into:
x = 30 - 5y
-5 = 30 - 5y
-35 = -5y
7 = y
.
Solution: (x,y) = (-5, 7)